# Complexity reduction of physical models: an equation-free approach by means of scaling

### v1

@@ 3.1 Optimal Scaling
Optimal Scaling (OS) proposed in [6] enables such a rational choice of $\theta_{1},\dots,\theta_{N_{\mathsf{x}}}\in(0,\infty)$, that ensures the deviation of magnitudes of $\lambda_{1},\dots,\lambda_{N_{\mathsf{l}}}$ (13) from unity to be minimal for any fixed choice of physical parameters $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$. This is advantageous for numerical computations as the floating-point numbers have the highest density in the interval $(0,1)$ and their density decreases when moving further away from this interval. So, it is always convenient to manipulate numerically the quantities with magnitude $\approx 1$, in order to avoid severe round-off errors and to improve the conditioning of the problem at hand.
In [6], such a methodology was first applied to liquid crystal and quantum models. The reported data demonstrate that OS ensures a strong reduction of the range of orders of magnitude assumed by involved quantities. Despite such a strong transformation, OS allows for an accurate estimation of characteristic features of considered systems. As an example, we showed that OS correctly describes the large-body limit of Landau-de Gennes model for liquid crystals [11]. Moreover, for Schrödinger Equation [12], we proved that the characteristic constants, provided by OS, follow the same trend as the often used atomic units.
Further, it was demonstrated that OS was also beneficial for the case of the Population Balance Model [8] for Latex Particles Morphology Formation [9, 10]. When compared with other scaling approaches [13, 14], it assured the smallest variation $\mathop{\mathrm{max}}_{i=1,\dots,N_{\mathsf{l}}}\lambda_{i}/\mathop{\mathrm{min}}_{i=1,\dots,N_{\mathsf{l}}}\lambda_{i}$ of coefficients (13) and allowed reducing round-off and numerical errors, as discussed in [6].
@@ 3.1.1 Formulation
Given any fixed choice of physical parameters $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$, OS proposed in [6] finds $\vec{\theta}=\vec{\theta}_{\mathrm{opt}}$ in such a way that the deviation of magnitudes of $\lambda_{1},\dots,\lambda_{N_{\mathsf{l}}}$ (13) from unity is minimal, i.e.,
@@
$\mbox{where}\quad\vec{\theta}\coloneqq(\theta_{1},\dots,\theta_{N_{\mathsf{x}}})\in(0,\infty)^{N_{\mathsf{x}}}\quad\mbox{and the constants}\quad\Theta_{1}=\cdots=\Theta_{N_{\mathsf{l}}}=0$ (14)
hold the targeted orders of magnitude, i.e., $\Theta_{i}=\log_{10}(1)=0$, $\forall i=1,\dots,N_{\mathsf{l}}$.
With $U=\{1,\dots,N_{\mathsf{l}}\}$, the function $C$ (14) is equivalent to $C_{U}$ in Theorem B.1 (Appendix B). Then, given Property A.1 (Appendix A) and any fixed choice of $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$, the points $\vec{\theta}_{\mathrm{opt}}$ of global minimum of $C\equiv C_{U}$ can be computed by means of (B.2), as follows from Theorem B.1 with $U=\{1,\dots,N_{\mathsf{l}}\}$. In this way, OS provides scaled coefficients $\lambda_{1},\dots,\lambda_{N_{\mathsf{l}}}$ (13) with magnitudes as close as possible to the targeted value $1$. In other words, OS attempts to make
@@
By its very design, OS finds the numerical values of the characteristic constants $\tilde{\theta}_{1},\dots,\tilde{\theta}_{N_{\mathsf{x}}}$ minimizing the Euclidean norm $C$ (14) of the vector $\left(\log_{10}(\lambda_{1}),\log_{10}(\lambda_{2}),\dots,\log_{10}(\lambda_{N_{\mathsf{l}}})\right)$. As discussed in [6], the Euclidean norm $C$ (14) is not the unique choice for the function to minimize. By considering different objective functions, we show in [6] that the performance of OS is not affected by the particular choice of the function in use. This observation can be motivated by the equivalence of norms on the finite dimensional space $(-\infty,\infty)^{N_{\mathsf{l}}}$. Nevertheless, the availability of an explicitly computable solution makes $C$ (14) the metric of choice for Optimal Scaling.
@@ 3.2 Optimal Scaling with Constraints
Because of the similar values achieved by $\lambda_{1},\dots,\lambda_{N_{\mathsf{l}}}$, the Optimal Scaling (OS) summarized above does not allow examining regimes of $\lambda_{1},\dots,\lambda_{N_{\mathsf{l}}}$ (13) in (12) which exhibit a large difference in magnitudes. The Optimal Scaling with Constraints (OSC) proposed in [7] addresses scenarios where (at least) one coefficient is much smaller than the others, i.e., most of $\lambda_{1},\dots,\lambda_{N_{\mathsf{l}}}$ in (12) are as close as possible to $1$, while (at least) one coefficient is $\ll 1$. The method quantifies those $\theta_{1},\dots,\theta_{N_{\mathsf{x}}}\in(0,\infty)$ that lead to such a regime.
As demonstrated in [7] on the example of the Population Balance Model [8] for Latex Particles Morphology Formation [9, 10], OSC can also determine a threshold on the values of physical parameters $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$ that allow dropping the terms relative to the $\lambda_{i}\ll 1$ in the considered equations without significantly changing their solutions, but dramatically reducing computational effort.
@@ 3.2.1 Formulation
Given any fixed choice of $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$ and following the procedure proposed in [7], it is possible to show that the vector $\vec{\theta}=\vec{\theta}^{\star}_{\mathrm{opt}}$ targeting the regime (16) of $\lambda_{1},\dots,\lambda_{N_{\mathsf{l}}}$ (13) can be found as
@@
and $C_{U_{\mathtt{s}}}$ accounts only for the coefficients $\lambda_{i}\approx 1$, as specified by $U_{\mathtt{s}}$ (16). By this means, the cost function $C_{U_{\mathtt{s}}}$ (17) measures the distance between the vector holding the orders of magnitude of coefficients $\lambda_{i}$, with $i\in U_{\mathtt{s}}$, and the corresponding vector of desired orders $\Theta_{i}=0$, with $i\in U_{\mathtt{s}}$. Assuming Property A.1 (Appendix A), provided $U=U_{\mathtt{s}}$ (16) and given any fixed $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$, the points $\vec{\theta}^{\star}_{\mathrm{opt}}$ of global minimum of $C_{U_{\mathtt{s}}}$ can be computed by means of (B.2), as follows from Theorem B.1 (Appendix B).
Then, the following features can be attained by the OSC methodology [7] and the computed $\vec{\theta}^{\star}_{\mathrm{opt}}$ (17). For any given $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$, the coefficients $\lambda_{i}$, with $i\in U_{\mathtt{s}}$ (16), result to be as close as possible to the targeted value $1$, i.e., OSC attempts to make
@@
with $\vec{\theta}^{\star}_{\mathrm{opt},j}$ being the $j$-th component of the vector $\vec{\theta}^{\star}_{\mathrm{opt}}$ (17). Hence, the scaled model (12) accounting only for the terms relative to $\lambda_{i}$, with $i\in U_{\mathtt{s}}$ (16), inherits the features of the OS approach [6] discussed in Section 3.1.
@@ 4 Generalized Optimal Scaling
Next, we intend to enrich the OSC method with an extra feature, namely the ability to reduce to its minimum the number of parameters employed in a model. With this addition, OSC can optionally control over-parametrization and should be beneficial for real-life models with a large number of physical parameters. This can be achieved by applying ideas and concepts relative to dimensional analysis and the Buckingham $\pi$ Theorem [15] in particular. From now on, we will refer to OSC with such a functionality as Generalized Optimal Scaling, or GOS.
@@ 4.1 Formulation
By requiring Properties A.1, A.2, A.3, A.4, A.5 and A.6 (Appendix A) and applying Theorems B.1, B.2, B.3, B.4, B.5, B.6 and B.7 (Appendix B), we formulate here the GOS methodology that assures all the features of OSC, while enables the minimization of the number of parameters employed.
Given Properties A.4, A.5 and A.6 of physical parameters $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$, Theorem B.3 assures that there exist only $N_{\mathsf{0}}$ independent dimensionless parameters $\pi_{1},\dots,\pi_{N_{\mathsf{0}}}\in(0,\infty)$ being a power-law monomial of $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$. This to say that any dimensionless power-law monomial of $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$ can be written in terms of $\pi_{1},\dots,\pi_{N_{\mathsf{0}}}$ only and, thus, $N_{\mathsf{0}}$ is the minimal number of dimensionless power-law monomials of $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$ one can consider.
By applying Properties A.1, A.2, A.3, A.5, one reveals that each coefficient $\lambda_{1},\dots,\lambda_{N_{\mathsf{l}}}$ (13) of model (12) is a dimensionless power-law monomial of $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$ and, thus, it can be written in terms of $\pi_{1},\dots,\pi_{N_{\mathsf{0}}}$ only, as stated by (B.20) in Theorem B.4 under Properties A.1, A.2, A.3, A.4, A.5 and A.6. This implies that (9) and (12) are equivalent to a model with governing coefficients given by $\pi_{1},\dots,\pi_{N_{\mathsf{0}}}$ only. For this reason, it is possible to reformulate the scaling approach in terms of the independent dimensionless parameters $\pi_{1},\dots,\pi_{N_{\mathsf{0}}}$, whose number $N_{\mathsf{0}}$ is minimal.
The formulation of the GOS methodology is based on replacing $\lambda_{i}$ in (17) with the expression (B.20) provided by Theorem B.4 under Properties A.1, A.2, A.3, A.4, A.5 and A.6. Then, the optimization problem (17) is considered in terms of the vector $\vec{\alpha}\in\mathbb{R}^{N_{\mathsf{l}}\,N_{\mathsf{0}}}$ (B.39) collecting the exponents $\alpha_{i,k}$ in (B.20). As dictated by Theorem B.6 when Properties A.1, A.2, A.3, A.4, A.5 and A.6 hold, such a vector $\vec{\alpha}\in\mathbb{R}^{N_{\mathsf{l}}\,N_{\mathsf{0}}}$ (B.39) must satisfy the linear system $\textbf{S}\,\vec{\alpha}=\vec{s}$ (B.40). Then, the optimization of $\vec{\alpha}$ must take into account such a constraint, as specified in what follows.
Given any fixed choice of $\pi_{1},\dots,\pi_{N_{\mathsf{0}}}\in(0,\infty)$, it is possible to target values of $\vec{\alpha}\in\mathbb{R}^{N_{\mathsf{l}}\,N_{\mathsf{0}}}$ (B.39) satisfying $\textbf{S}\,\vec{\alpha}=\vec{s}$ (B.40) and minimizing the deviation $C_{U_{\mathtt{s}}}:\mathbb{R}^{N_{\mathsf{l}}\,N_{\mathsf{0}}}\to[0,\infty)$ from unity of the magnitudes of $\lambda_{i}$, with $i\in U_{\mathtt{s}}\coloneqq\{1,\dots,N_{\mathsf{l}}\}\setminus\{\mathtt{s}\}$ for a fixed $\mathtt{s}\in\{0,1,\dots,N_{\mathsf{l}}\}$:
@@
Given $U=U_{\mathtt{s}}$, let $\boldsymbol{\Sigma}=\textbf{S}$ and $\vec{\sigma}=\vec{s}$ satisfy (B.50). Then, Theorem B.7 allows computing $\vec{\alpha}_{\mathrm{opt}}$ (20) by means of (B.49). If $\boldsymbol{\Sigma}=\textbf{S}$ and $\vec{\sigma}=\vec{s}$ do not fulfill (B.50), but such an assumption is accomplished by the row echelon form of $\textbf{S}\,\vec{\alpha}=\vec{s}$, it is possible to assign $\boldsymbol{\Sigma}$ and $\vec{\sigma}$ as the corresponding matrices of such a reduced system.
@@
Employed Assumptions Property A.1 Properties A.1, A.2, A.3, A.4, A.5 and A.6
@@
Procedure Given any fixed choice of physical parameters $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$, finds $\vec{\theta}\coloneqq(\theta_{1},\dots,\theta_{N_{\mathsf{x}}})\in(0,\infty)^{N_{\mathsf{x}}}$ that minimizes the deviation $C_{U_{\mathtt{s}}}$ (17) from unity of the magnitudes of $\lambda_{i}$, $\forall i\in U_{\mathtt{s}}\coloneqq\{1,\dots,N_{\mathsf{l}}\}\setminus\{\mathtt{s}\}$ Given any fixed choice of $\pi_{1},\dots,\pi_{N_{\mathsf{0}}}\in(0,\infty)$, finds $\vec{\alpha}$ (B.39) that satisfies (B.40) and minimizes the deviation $C_{U_{\mathtt{s}}}$ (20) from unity of the magnitudes of $\lambda_{i}$, $\forall i\in U_{\mathtt{s}}\coloneqq\{1,\dots,N_{\mathsf{l}}\}\setminus\{\mathtt{s}\}$
Realization The points $\vec{\theta}$ of global minimum of $C_{U_{\mathtt{s}}}$ (17) can be computed by means of (B.2) provided $U=U_{\mathtt{s}}$, as follows from Theorem B.1 With $U=U_{\mathtt{s}}$, $\boldsymbol{\Sigma}=\textbf{S}$, $\vec{\sigma}=\vec{s}$ (B.40) and assuming (B.50), Theorem B.7 allows computing $\vec{\alpha}=\vec{\alpha}_{\mathrm{opt}}$ (20) by means of (B.49)
Plugs $\vec{\alpha}=\vec{\alpha}_{\mathrm{opt}}$ (20) into (B.38), then solves for $\vec{t}$ (B.24) the system (B.25) and uses the found $\vec{t}$ to compute $\tilde{\theta}_{1},\dots,\tilde{\theta}_{N_{\mathsf{x}}}$ as in (A.3)
Table 1: Summary and features of Optimal Scaling with Constraints [7] and Generalized Optimal Scaling (GOS). The Optimal Scaling [6] is equivalent to GOS with $\mathtt{s}=0$.

### v2

@@ 3.2 Optimal Scaling
Optimal Scaling (OS) proposed in [RUSCONI2019106944] enables such a rational choice of $\theta_{1},\dots,\theta_{N_{\mathsf{x}}}\in(0,\infty)$, that ensures the deviation of magnitudes of $\lambda_{1},\dots,\lambda_{N_{\mathsf{l}}}$ (13) from unity to be minimal for any fixed choice of physical parameters $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$. This is advantageous for numerical computations as the floating-point numbers have the highest density in the interval $(0,1)$ and their density decreases when moving further away from this interval. So, it is always convenient to manipulate numerically the quantities with magnitude $\approx 1$, in order to avoid severe round-off errors and to improve the conditioning of the problem at hand.
In [RUSCONI2019106944], such a methodology was first applied to liquid crystal and quantum models. The reported data demonstrate that OS ensures a strong reduction of the range of orders of magnitude assumed by involved quantities. Despite such a strong transformation, OS allows for an accurate estimation of characteristic features of considered systems. As an example, we showed that OS correctly describes the large-body limit of Landau-de Gennes model for liquid crystals [Gartland2018]. Moreover, for Schrödinger Equation [PhysRevA.53.2135], we proved that the characteristic constants, provided by OS, follow the same trend as the often used atomic units.
Further, it was demonstrated that OS was also beneficial for the case of the Population Balance Model [Ramkrishna2000] for Latex Particles Morphology Formation [DDPM_2016, PhDThesis_Rusconi_PMCQS]. When compared with other scaling approaches [Holmes2009_ND, Scaling_Langtangen], it assured the smallest variation $\mathop{\mathrm{max}}_{i=1,\dots,N_{\mathsf{l}}}\lambda_{i}/\mathop{\mathrm{min}}_{i=1,\dots,N_{\mathsf{l}}}\lambda_{i}$ of coefficients (13) and allowed reducing round-off and numerical errors, as discussed in [RUSCONI2019106944].
@@ 3.2.1 Formulation
Given any fixed choice of physical parameters $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$, OS proposed in [RUSCONI2019106944] finds $\vec{\theta}=\vec{\theta}_{\mathrm{opt}}$ in such a way that the deviation of magnitudes of $\lambda_{1},\dots,\lambda_{N_{\mathsf{l}}}$ (13) from unity is minimal, i.e.,
@@
$\mbox{where}\quad\vec{\theta}\coloneqq(\theta_{1},\dots,\theta_{N_{\mathsf{x}}})\in(0,\infty)^{N_{\mathsf{x}}}\quad\mbox{and the constants}\quad\Theta_{1}=\cdots=\Theta_{N_{\mathsf{l}}}=0.$ (33)
In (33), the order of magnitude of $\lambda_{i}$ is computed as $\log_{10}(\lambda_{i})$, while $\Theta_{i}=0$ to target the desired order $\log_{10}(1)=0$, $\forall i=1,\dots,N_{\mathsf{l}}$.
With $U=\{1,\dots,N_{\mathsf{l}}\}$, the function $C$ (33) is equivalent to $C_{U}$ in Theorem 6 (Section 3.1.2). Then, given Assumption˜1 (Section 3.1) and any fixed choice of $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$, the points $\vec{\theta}_{\mathrm{opt}}$ of global minimum of $C\equiv C_{U}$ can be computed by means of (28), as follows from Theorem 6 with $U=\{1,\dots,N_{\mathsf{l}}\}$. In this way, OS provides scaled coefficients $\lambda_{1},\dots,\lambda_{N_{\mathsf{l}}}$ (13) with magnitudes as close as possible to the targeted value $1$. In other words, OS attempts to make
@@
By its very design, OS finds the numerical values of the characteristic constants $\tilde{\theta}_{1},\dots,\tilde{\theta}_{N_{\mathsf{x}}}$ minimizing the Euclidean norm $C$ (33) of the vector $\left(\log_{10}(\lambda_{1}),\log_{10}(\lambda_{2}),\dots,\log_{10}(\lambda_{N_{\mathsf{l}}})\right)$. As discussed in [RUSCONI2019106944], the Euclidean norm $C$ (33) is not the unique choice for the function to minimize. By considering different objective functions, we show in [RUSCONI2019106944] that the performance of OS is not affected by the particular choice of the function in use. This observation can be motivated by the equivalence of norms on the finite dimensional space $(-\infty,\infty)^{N_{\mathsf{l}}}$. Nevertheless, the availability of an explicitly computable solution makes $C$ (33) the metric of choice for Optimal Scaling.
@@ 3.3 Optimal Scaling with Constraints
Because of the similar values achieved by $\lambda_{1},\dots,\lambda_{N_{\mathsf{l}}}$, the Optimal Scaling (OS) summarized above does not allow examining regimes of $\lambda_{1},\dots,\lambda_{N_{\mathsf{l}}}$ (13) in (12) which exhibit a large difference in magnitudes. The Optimal Scaling with Constraints (OSC) proposed in [RUSCONI2023127756] addresses scenarios where (at least) one coefficient is much smaller than the others, i.e., most of $\lambda_{1},\dots,\lambda_{N_{\mathsf{l}}}$ in (12) are as close as possible to $1$, while (at least) one coefficient is $\ll 1$. The method quantifies those $\theta_{1},\dots,\theta_{N_{\mathsf{x}}}\in(0,\infty)$ that lead to such a regime.
As demonstrated in [RUSCONI2023127756] on the example of the Population Balance Model [Ramkrishna2000] for Latex Particles Morphology Formation [DDPM_2016, PhDThesis_Rusconi_PMCQS], OSC can also determine a threshold on the values of physical parameters $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$ that allow dropping the terms relative to the $\lambda_{i}\ll 1$ in the considered equations without significantly changing their solutions, but dramatically reducing computational effort.
@@ 3.3.1 Formulation
Given any fixed choice of $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$ and following the procedure proposed in [RUSCONI2023127756], it is possible to show that the vector $\vec{\theta}=\vec{\theta}^{\star}_{\mathrm{opt}}$ targeting the regime (35) of $\lambda_{1},\dots,\lambda_{N_{\mathsf{l}}}$ (13) can be found as
@@
and $C_{U_{\mathtt{s}}}$ accounts only for the coefficients $\lambda_{i}\approx 1$, as specified by $U_{\mathtt{s}}$ (35). Moreover, we remark that (36) quantifies the order of magnitude of $\lambda_{i}$ as $\log_{10}(\lambda_{i})$, while $\Theta_{i}=0$ to target the desired order $\log_{10}(1)=0$, $\forall i=1,\dots,N_{\mathsf{l}}$. By this means, the cost function $C_{U_{\mathtt{s}}}$ (36) measures the distance between the vector holding the orders of magnitude of coefficients $\lambda_{i}$, with $i\in U_{\mathtt{s}}$, and the corresponding vector of desired orders $\Theta_{i}=0$, with $i\in U_{\mathtt{s}}$. Made Assumption˜1 (Section 3.1), provided $U=U_{\mathtt{s}}$ (35) and given any fixed $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$, the points $\vec{\theta}^{\star}_{\mathrm{opt}}$ of global minimum of $C_{U_{\mathtt{s}}}$ can be computed by means of (28), as follows from Theorem 6 (Section 3.1.2).
Then, the following features can be attained by the OSC methodology [RUSCONI2023127756] and the computed $\vec{\theta}^{\star}_{\mathrm{opt}}$ (36). For any given $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$, the coefficients $\lambda_{i}$, with $i\in U_{\mathtt{s}}$ (35), result to be as close as possible to the targeted value $1$, i.e., OSC attempts to make
@@
with $\vec{\theta}^{\star}_{\mathrm{opt},j}$ being the $j$-th component of the vector $\vec{\theta}^{\star}_{\mathrm{opt}}$ (36). Hence, the scaled model (12) accounting only for the terms relative to $\lambda_{i}$, with $i\in U_{\mathtt{s}}$ (35), inherits the features of the OS approach [RUSCONI2019106944] discussed in Section 3.2.
@@ 4 Generalized Optimal Scaling
Next, we intend to enrich the OSC method with an extra feature, namely the ability to reduce to its minimum the number of parameters employed in a model. With this addition, OSC can optionally control over-parametrization and should be beneficial for real-life models with a large number of physical parameters. This can be achieved by applying ideas and concepts relative to dimensional analysis and the Buckingham $\pi$ Theorem [PhysRev.4.345] in particular. From now on, we will refer to OSC with such a functionality as Generalized Optimal Scaling, or GOS.
@@ 4.1 Formulation
By requiring Assumptions˜1, 2, 3, 4, 5 and 6 (Section 3.1) and applying Theorems˜1, 2, 3, 4, 5 and 7 (Section 3.1.1 and Section 3.1.2), we formulate here the GOS methodology that assures all the features of OSC, while enables the minimization of the number of parameters employed.
Given Assumptions˜4, 5 and 6 for physical parameters $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$, Theorem 1 and Theorem 2 assure that there exist only $N_{\mathsf{0}}$ independent dimensionless parameters $\pi_{1},\dots,\pi_{N_{\mathsf{0}}}\in(0,\infty)$ being a power-law monomial of $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$. This to say that any dimensionless power-law monomial of $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$ can be written in terms of $\pi_{1},\dots,\pi_{N_{\mathsf{0}}}$ only and, thus, $N_{\mathsf{0}}$ is the minimal number of dimensionless power-law monomials of $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$ one can consider.
By invoking Assumptions 1, 2 and 3, one reveals that each coefficient $\lambda_{1},\dots,\lambda_{N_{\mathsf{l}}}$ (13) of model (12) is a dimensionless power-law monomial of $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$ and, thus, it can be written in terms of $\pi_{1},\dots,\pi_{N_{\mathsf{0}}}$ only, as stated by (22) in Theorem 3 under Assumptions˜1, 2, 3, 4, 5 and 6. This implies that (9) and (12) are equivalent to a model with governing coefficients given by $\pi_{1},\dots,\pi_{N_{\mathsf{0}}}$ only. For this reason, it is possible to reformulate the scaling approach in terms of the independent dimensionless parameters $\pi_{1},\dots,\pi_{N_{\mathsf{0}}}$, whose number $N_{\mathsf{0}}$ is minimal.
The formulation of the GOS methodology is based on replacing $\lambda_{i}$ in (36) with the expression (22) provided by Theorem 3 under Assumptions˜1, 2, 3, 4, 5 and 6. Then, the optimization problem (36) is considered in terms of the vector $\vec{\alpha}\in\mathbb{R}^{N_{\mathsf{l}}\,N_{\mathsf{0}}}$ (25) collecting the exponents $\alpha_{i,k}$ in (22). As dictated by Theorem 4 and Theorem 5 when Assumptions˜1, 2, 3, 4, 5 and 6 hold, such a vector $\vec{\alpha}\in\mathbb{R}^{N_{\mathsf{l}}\,N_{\mathsf{0}}}$ (25) must satisfy the linear system $\textbf{S}\,\vec{\alpha}=\vec{s}$ (26). Then, the optimization of $\vec{\alpha}$ must take into account such a constraint, as specified in what follows.
Given any fixed choice of $\pi_{1},\dots,\pi_{N_{\mathsf{0}}}\in(0,\infty)$, it is possible to target values of $\vec{\alpha}\in\mathbb{R}^{N_{\mathsf{l}}\,N_{\mathsf{0}}}$ (25) satisfying $\textbf{S}\,\vec{\alpha}=\vec{s}$ (26) and minimizing the deviation $C_{U_{\mathtt{s}}}:\mathbb{R}^{N_{\mathsf{l}}\,N_{\mathsf{0}}}\to[0,\infty)$ from unity of the magnitudes of $\lambda_{i}$, with $i\in U_{\mathtt{s}}\coloneqq\{1,\dots,N_{\mathsf{l}}\}\setminus\{\mathtt{s}\}$ for a fixed $\mathtt{s}\in\{0,1,\dots,N_{\mathsf{l}}\}$:
@@
Given $U=U_{\mathtt{s}}$, let $\boldsymbol{\Sigma}=\textbf{S}$ and $\vec{\sigma}=\vec{s}$ satisfy (32). Then, Theorem 7 allows computing $\vec{\alpha}_{\mathrm{opt}}$ (39) by means of (31). If $\boldsymbol{\Sigma}=\textbf{S}$ and $\vec{\sigma}=\vec{s}$ do not fulfill (32), but such an assumption is accomplished by the row echelon form of $\textbf{S}\,\vec{\alpha}=\vec{s}$, it is possible to assign $\boldsymbol{\Sigma}$ and $\vec{\sigma}$ as the corresponding matrices of such a reduced system.
@@
Employed Assumptions Assumption˜1 Assumptions˜1, 2, 3, 4, 5 and 6
@@
Procedure Given any fixed choice of physical parameters $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$, finds $\vec{\theta}\coloneqq(\theta_{1},\dots,\theta_{N_{\mathsf{x}}})\in(0,\infty)^{N_{\mathsf{x}}}$ that minimizes the deviation $C_{U_{\mathtt{s}}}$ (36) from unity of the magnitudes of $\lambda_{i}$, $\forall i\in U_{\mathtt{s}}\coloneqq\{1,\dots,N_{\mathsf{l}}\}\setminus\{\mathtt{s}\}$ Given any fixed choice of $\pi_{1},\dots,\pi_{N_{\mathsf{0}}}\in(0,\infty)$, finds $\vec{\alpha}$ (25) that satisfies (26) and minimizes the deviation $C_{U_{\mathtt{s}}}$ (39) from unity of the magnitudes of $\lambda_{i}$, $\forall i\in U_{\mathtt{s}}\coloneqq\{1,\dots,N_{\mathsf{l}}\}\setminus\{\mathtt{s}\}$
Realization The points $\vec{\theta}$ of global minimum of $C_{U_{\mathtt{s}}}$ (36) can be computed by means of (28) provided $U=U_{\mathtt{s}}$, as follows from Theorem 6 With $U=U_{\mathtt{s}}$, $\boldsymbol{\Sigma}=\textbf{S}$, $\vec{\sigma}=\vec{s}$ (26) and assuming (32), Theorem 7 allows computing $\vec{\alpha}=\vec{\alpha}_{\mathrm{opt}}$ (39) by means of (31)
Plugs $\vec{\alpha}=\vec{\alpha}_{\mathrm{opt}}$ (39) into (D.13), then solves for $\vec{t}$ (23) the system (24) and uses the found $\vec{t}$ to compute $\tilde{\theta}_{1},\dots,\tilde{\theta}_{N_{\mathsf{x}}}$ as in (16)
Table 1: Summary and features of Optimal Scaling with Constraints [RUSCONI2023127756] and Generalized Optimal Scaling (GOS). The Optimal Scaling [RUSCONI2019106944] is equivalent to GOS with $\mathtt{s}=0$.